The vertex of the parabola below is at the point (3, 2) and the point (4, 6) is on the parabola. What is the equation of the parabola?
a. x = 6(y - 3)^2 + 2
b. y = 4(x - 3)^2 + 2
c. y = 2(x - 2)^2 + 3
d. y = 4(x + 3)^2 - 2
..Please !!!

Answer by lwsshak3(6509)

The vertex of the parabola below is at the point (3, 2) and the point (4, 6) is on the parabola. What is the equation of the parabola?
a. x = 6(y - 3)^2 + 2
b. y = 4(x - 3)^2 + 2
c. y = 2(x - 2)^2 + 3
d. y = 4(x + 3)^2 - 2
**
Standard form of equation for a parabola: y=A(x-h)^2+k, (h,k)=(x,y) coordinates of vertex, A is a multiplier that affects the slope or steepness of the curve.
y=A(x-h)^2+k
plug in vertex coordinates (3,2)
y=A(x-3)^2+2
plug in coordinates of given point (4,6)
6=A(4-3)^2+2
6=A(1)^2+2
6=A+2
A=4
Equation:
y=4(x-3)^2+2
b. is the answer